# Minimising surface-code failures using a color-code decoder

### v2

@@ Abstract
The development of practical, high-performance decoding algorithms reduces the resource cost of fault-tolerant quantum computing. Here we propose a decoder for the surface code that finds low-weight correction operators for errors produced by the depolarising noise model. The decoder is obtained by mapping the syndrome of the surface code onto that of the color code, thereby allowing us to adopt more sophisticated color-code decoding algorithms. Analytical arguments and exhaustive testing show that the resulting decoder can find a least-weight correction for all weight $d/2$ depolarising errors for even code distance $d$. This improves the logical error rate by an exponential factor $O(2^{d/2})$ compared with decoders that treat bit-flip and dephasing errors separately. We demonstrate this improvement with analytical arguments and supporting numerical simulations at low error rates. Of independent interest, we also demonstrate an exponential improvement in logical error rate for our decoder used to correct independent and identically distributed bit-flip errors affecting the color code compared with more conventional color-code decoding algorithms.
@@ 1 Introduction
Topological codes, such as the surface code [15, 16, 7, 17] and color code [18, 19, 20, 21], are among the most promising quantum error-correcting codes to be realised with quantum technology that is now under development [22, 23, 24, 25, 26, 27]. This is due to their layout which can be realised with a planar array of qubits with only nearest-neighbour interactions, and their demonstrated high threshold error rates, such that they can function under significant laboratory noise. A well-established class of decoders for the surface code [7, 28, 29, 17] is based on the minimum-weight perfect-matching algorithm (MWPM) [30, 31]. We colloquially refer to these decoders as matching decoders. Matching decoders give rise to high thresholds for topological codes, and have been demonstrated to be highly versatile in that they can be adapted to correct for different noise models [32, 33, 34, 35, 36, 37, 38, 39, 40], as well as different codes [41, 42, 43], including the color code [44, 45, 46, 47, 48, 49] and surface code variants [50, 51, 52, 53, 54]. See the perspective article, Ref. [55] for a discussion.
@@
A matching decoder exploits the structure of the code to identify errors which likely caused the error syndrome. Specifically, we take subsets of the syndrome data and represent them on a lattice, such that this syndrome data respects certain symmetries of the code [15, 41, 55]. These symmetries are characterised by the fact that errors produce syndrome violations, commonly known as defects, in spatially local pairs over the lattice. Given such a lattice, the MWPM algorithm can be used to pair spatially local defects. In turn, this pairing can be interpreted to identify an error that is likely to have given rise to the syndrome, such that a correction that recovers the encoded state can be determined.
@@
The performance of a matching decoder is determined by the choice of syndrome lattice. The surface code has consistently demonstrated high thresholds by concentrating on syndrome lattices which correct for bit-flip errors and dephasing errors separately. However, without additional decoding steps [56, 45, 57], such decoders cannot identify correlations between these two error types that occur in common error models such as depolarising noise. See for example Refs. [8, 56] where candid discussions are presented on the limitations of matching decoder variations correcting for depolarising noise. In this work, we introduce a syndrome lattice for the surface code that accounts for correlations between bit-flip and dephasing errors, which we obtain by exploiting a correspondence between the surface code and the color code, commonly known as an unfolding map [58, 59, 60, 61]. Through this mapping, we can adopt color-code decoding algorithms to correct noise on the surface code. Indeed, a syndrome lattice for our decoder naturally ensues by adapting the decoding methods introduced in Ref. [48].
@@
We find that our unified decoder determines a least-weight correction for all weight $d/2$ depolarising noise errors for surface codes with even distance $d$. As such, in the limit of low error rate, we obtain optimal logical failure rates. In contrast, conventional matching decoders that do not account for correlated errors have a logical failure rate $O(2^{d/2})$ higher than the decoder we present. We adopt a number of methods to analyse the performance of our decoder at low error rates. We first evaluate the number of weight $d/2$ errors that should lead to a logical failure for both our matching decoder, as well as a conventional matching decoder. We verify our expressions with exhaustive searching of weight $d/2$ errors. In addition, we measure the logical failure rate of our decoder using the splitting method [62, 63]. We find the numerical results we obtain to be consistent with our analytic results in the limit of vanishing physical error rate, thereby verifying the performance of our decoder in comparison to conventional decoders. We also show that the threshold of our decoder is comparable to more conventional decoding methods. Finally, we investigate the performance of our scheme to decode the color code undergoing bit-flip noise. We find our analysis sheds light on the role of entropy in quantum error correction with the matching decoders that are central to our study [64, 65, 66].
@@ 2.1 The stabilizer formalism
Quantum-error correcting codes are designed to protect logical information that is encoded in a subspace of a larger physical Hilbert space. This subspace is called the code space. A large class of QEC codes are readily described using the stabilizer formalism [67], whereby a QEC code is defined by an Abelian subgroup of the Pauli group $\mathcal{P}$, known as the stabilizer group $\mathcal{S}\subset\mathcal{P}$. Up to phases, the Pauli group on $n$ qubits is generated by $n$-qubit Pauli matrices; $X_{q}$ and $Z_{q}$ respectively denote, the standard Pauli-X and Pauli-Z operator acting on qubit $q$. The code space of a stabilizer code is the +$1$ eigenspace of all of its elements, namely $s\ket{\psi}=\ket{\psi}$ $\forall$ $s\in\mathcal{S}$, where the code space is spanned by eigenstates |$\ket{\psi}$.
@@
A correction $C$ can be found and applied to return the code to the ground subspace such that $CE\in\mathcal{S}$, and recover the encoded logical state |$\ket{\psi}$. We obtain $C$ using a decoding algorithm that is designed to determine a highly probable error giving rise to a measured syndrome, given a known error model where we assume errors are introduced at a low rate. In general, an $\llbracket n,k,d\rrbracket$ quantum code is characterised by the number of physical qubits $n$, the number of logical qubits $k$ that they encode, and a code distance $d$ given by the minimum-weight of an error configuration triggering a logical error.
@@ 2.2 The surface code
In this work, we propose and implement a matching decoder to correct for the surface code undergoing an independent and identically distributed depolarising noise model. This error model is such that each qubit experiences either a Pauli-$X$, -$Y$, or -$Z$ error with probability $p/3$, or no error with probability $(1-p)$. We indicate their corresponding syndromes in Fig. 1.
@@ 2.3 The color code
The color code is defined on a three-colourable two-dimensional lattice. We will focus on the regular square-octagon color code with square boundaries, see Fig. 2. Physical qubits are located on the vertices of the lattice, $v$, and each face $f$ supports two stabilizer operators $S_{f}^{P}=\prod_{v\in E_{f}}P_{e}$ where $P\in\{X,Z\}$ and $E_{f}$ denotes the qubits on the boundaries of face $f$.
We define a set of colour labels red, green and blue in bold $\mathcal{C}=\{\textbf{r},\textbf{g},\textbf{b}\}$, and each face of the code is assigned a label $\textbf{u}\in\mathcal{C}$ such that no two faces of same colour touch.
@@
We will consider a correspondence between depolarising noise on the surface and two-qubit bit-flip events on the color code. These give rise to error strings $E=\prod_{v\in E_{v}}X_{v_{1}}X_{v_{2}}$, where $E_{v}$ is a subset of red stabilizers on the color code affected by two bit-flip errors on qubits $v_{1}$ and $v_{2}$. We note that a defect is attributed the colour from $\mathcal{C}$ of the violated stabilizer it lies on, and discuss the decoding of the color code in Sec. 3.
@@ 3 Materialised symmetries and decoding
In this section, we review the concept of materialised symmetries for topological quantum codes introduced in Refs. [15, 41, 48, 55]. These symmetries give rise to a parity conservation law among syndrome defects that we exploit to make a matching decoder for the square-boundary color code. In Ref. [48], a global symmetry termed the unified lattice was derived for the triangular color-code with a single red, blue and green boundary. Notably, the unified lattice presented in Ref. [48] had the topology of a Möbius strip. Motivating the utility of a unified lattice, the work showed that both the threshold, and the logical failure rates were improved for the triangular color code. As such, we introduce a unified decoder for the color code presented in Sec. 2.3. In addition to introducing the unified lattice decoder, we will also review the matching decoder on the restricted lattice [44, 45, 46, 47, 49]. In later sections we will compare the restricted-lattice decoder to the unified-lattice decoder.
@@ 3.1 Symmetries of the color code
We find the symmetries of the color code by considering subsets of its stabilizer generators with specific colours [45, 48]. Our goal is to derive a single ‘unifying’ symmetry for the color code with boundaries. We also investigate other color code decoders based on restricted lattices. Interestingly, we find that the square-boundary color code has a unified lattice with a topology that is distinct from that in Ref. [48], and is given by the manifold in Fig. 3. In this subsection we show how to derive this unified lattice for the square boundary color code, which we will adopt in our matching decoder implementation.
@@
We show the restricted lattices $\mathcal{R}_{\mathbf{b}},\mathcal{R}_{\mathbf{r}},$ and $\mathcal{R}_{\mathbf{g}}$ in Fig. 4(a). Errors in the bulk of the color code preserve defect parity on any single restricted lattice. We show examples of bulk errors on the restricted lattices in Fig. 4(a), where the bottom left error on the lattice flips a single qubit and creates a single defect of each colour, and the top right error flips two qubits, and creates two green defects.
@@
where we take the product over all faces of the restricted lattice, i.e. those not coloured $\mathbf{u}$. This operator is highlighted for each restricted lattice in Fig. 4(b). We note that the boundary operator $b_{\mathbf{u}}$ is supported on all boundaries and corners that are not $\mathbf{u}$-coloured.
@@
3.2 Matching decoders code symmetries
@@
The unified-lattice decoder is obtained from the global symmetry of the color code described in the previous subsection, such that errors give rise to pairs of defects. Given a collection of defects on the unified lattice, the matching-based decoding problem reduces to finding a pairing between these defects that is created by an underlying error with maximal probability. Towards this goal, a matching between a pair of defects is assigned a weight inversely dependent on the probability of errors on all the qubits along the shortest path connecting the two. As a result, decoding is equivalent to finding the perfect matching of minimum weight on this weighted graph.
@@
Given a matching, we now lay out the procedure to obtain the relevant correction to the logical qubit subjected to $X$ errors on the underlying physical qubits. Note that the following procedure is valid for both the restriction and unified decoders. We first apply the Pauli-$Z$ corrections to qubits along the shortest paths between the pairs of defects determined by the matching algorithm. This returns the logical qubit to the code space. During this process, we track the parity of a logical operator, here the vertical $Z$ logical depicted in Fig. 5. Consider the case where the decoding algorithm has determined that pairs of defects are matched across this logical an odd (even) number of times. We successfully return the encoded qubit to the correct logical state if there are an odd (even) number of Pauli errors on the physical qubits along the logical, a method discussed in more detail in Ref. [55, 48]. In Fig. 5, if the decoder finds the yellow matching, the corresponding correction, together with the initial error, applies $\overline{X}$ to the encoded qubit; the grey matching returns the lattice to the correct state. Henceforth, we will refer to the MWPM-based decoder applied to syndrome data on the restricted (unified) lattice as the restricted (unified) decoder.
4 Mapping from surface to color code
@@
We have introduced the surface code and the color code in the previous section, and proposed a decoding procedure for bit-flip errors acting on the color code by matching on either the restricted lattice or the unified lattice. We now describe a local mapping from the surface code onto the square-octagonal color code, and introduce a mapping from depolarising errors on the former to correlated bit-flip errors on the latter. We exploit this mapping to adopt the restricted-lattice and unified-lattice color-code decoder for the surface code undergoing depolarising noise.
@@
Let us begin by stating the equivalence between the different lattice elements under this mapping, by considering the surface code displayed with a blue lattice in Fig. 6(a). First, we map the vertex and face operators of this blue surface code onto octagonal face operators of the color code. Specifically, each surface code star operator corresponds to a blue octagon $S_{f}$ of the color code and, likewise, each face operator of the surface code corresponds to a green octagonal face operator in the color-code picture $S_{f}$. This is shown in Fig. 6(a), where the star and face operators of the blue surface code align with their corresponding color code stabilizers exactly, thereby illustrating this mapping. We also show the surface code displayed with a green lattice, which represents a rotation of the blue counterpart, and where a similar argument holds.
Now that we have identified stars and plaquettes of each surface code with octagonal face operators of the color code, we must choose errors in the color-code picture that give rise to equivalent syndromes in the surface code picture. A Pauli-$Z$ error on the surface code will violate its two adjacent vertex operators, as shown in Fig. 1. In the color code picture, this error will violate the corresponding blue octagonal operators as shown in Fig. 6(b)(i) where we indicate this $Z$-flip on the blue surface code edge. To obtain this violation in the color code picture, we add two horizontal bit-flip errors onto the single red face of the color code on which the edge lies such that its two adjacent blue octagons support a syndrome, see Fig. 6(a). Similarly, a Pauli-$X$ error in the surface code picture will violate two adjacent face operators, which correspond to green octagons in the color code picture in Fig. 6(b)(ii). We find that the error in the color code picture that violates these appropriate color code octagons is a pair of bit-flip errors lying vertically on a single red square-shaped face operator shown in the corresponding diagram in Fig. 6(a).
@@
We have now proposed an error model that violates pairs of either blue or green face operators according to the mapping from the surface code undergoing a depolarising noise model. As we have alluded, single qubit errors in the surface code picture correspond to two-qubit errors supported on individual red faces of the color code. Indeed, we see that the edges in the surface code picture each have an underlying red face operator in the color-code picture, where the red square inherits an orientation from the underlying edge. We therefore identify qubits in the surface code picture with red faces in the color code picture. In what follows, it will be helpful to keep this identification in mind as we describe the mapping more rigorously. We note that this mapping is non-trivial in the sense that the commutation relations of the errors are not preserved. Nevertheless, we find a valid decoder for the surface code with the equivalence we have proposed.
@@
We then express the stabilizers of the blue and green surface codes respectively, using the two logical operators of the inner code $\{\overline{X}_{1},\overline{Z}_{1}\}$ and $\{\overline{X}_{2},\overline{Z}_{2}\}$, where the first(second) qubit encodes the blue(green) copy of the surface code, as shown in Fig. 6(d) (Fig. 6(e)). We emphasize that that the orientation of the edge qubit of the surface code dictates the orientation of the encoding of the $\llbracket 4,2,2\rrbracket$ code. See Fig. 6(c). Indeed, a red square associated with a horizontal (vertical) edge of the blue (green) surface code is encoded as per Fig. 6(c), but when traversed by a vertical edge, this encoding must be rotated by 90° to preserve our mapping.
@@
Finally, since the surface code on the green lattice is a duplicate of its blue counterpart up to a Hadamard rotation, we identify certain logical operators of the inner code to complete our mapping. Indeed, under the duplication, with the Hadamard rotation, we have the following relations between pairs of logical operators of the inner code, namely $\overline{X}_{1}\equiv\overline{Z}_{2}$ and $\overline{Z}_{1}\equiv\overline{X}_{2}$. This ensures the syndrome correspondence in Fig. 6(a) such that all depolarising errors on the surface code now have a one-to-one correspondence with a commuting two-qubit bit-flip error model. Most importantly, this enables us to use a decoder for a CSS code [3] to correct after depolarising noise on the surface code, represented in the color code picture.

### v3

# Minimising surface code failures using a color code decoder

@@ Abstract
The development of practical, high-performance decoding algorithms reduces the resource cost of fault-tolerant quantum computing. Here we propose a decoder for the surface code that finds low-weight correction operators for errors produced by the depolarising noise model. The decoder is obtained by mapping the syndrome of the surface code onto that of the color code, thereby allowing us to adopt more sophisticated graph-based decoding protocols used for the color code. Analytical arguments and exhaustive testing show that the resulting decoder can find a least-weight correction for all weight $d/2$ depolarising errors for even code distance $d$. This improves the logical error rate by an exponential factor $O(2^{d/2})$ compared with decoders that treat bit-flip and dephasing errors separately. We demonstrate this improvement with analytical arguments and supporting numerical simulations at low error rates. Of independent interest, we also demonstrate an exponential improvement in logical error rate for our decoder used to correct independent and identically distributed bit-flip errors affecting the color code compared with more conventional color code decoding algorithms.
@@ 1 Introduction
Topological codes, such as the surface code [15, 16, 7, 17] and color code [18, 19, 20, 21], are among the most promising quantum error-correcting codes to be realised with quantum technology that is now under development [22, 23, 24, 25, 26, 27]. This is due to their layout which can be realised with a planar array of qubits with only nearest-neighbour interactions, and their demonstrated high threshold error rates, such that they can function under significant laboratory noise. A well-established class of decoders for the surface code [7, 28, 29, 17] is based on the minimum-weight perfect-matching algorithm (MWPM) [30, 7, 31]. We colloquially refer to these decoders as matching decoders. Matching decoders give rise to high thresholds for topological codes, and have been demonstrated to be highly versatile in that they can be adapted to correct for different noise models [32, 33, 34, 35, 36, 37, 38, 39, 40], as well as different codes [41, 42, 43], including the color code [44, 45, 46, 47, 48, 49] and surface code variants [50, 51, 52, 53, 54]. See the perspective article, Ref. [55] for a discussion.
@@
A matching decoder exploits the structure of the code to identify errors which likely caused the error syndrome. Specifically, we take subsets of the syndrome data and represent them on a lattice, such that this syndrome data respects certain symmetries of the code [15, 41, 55]. These symmetries are characterised by the fact that errors produce syndrome violations, commonly known as defects, in pairs over the lattice. Given such a lattice, the MWPM algorithm can be used to pair spatially local defects. In turn, this pairing can be interpreted to identify an error that is likely to have given rise to the syndrome, such that a correction that recovers the encoded state can be determined.
@@
The performance of a matching decoder is determined by the choice of syndrome lattice. The surface code has consistently demonstrated high thresholds by concentrating on syndrome lattices which correct for bit-flip errors and dephasing errors separately. However, without additional decoding steps [56, 45, 57], such decoders cannot identify correlations between these two error types that occur in common error models such as depolarising noise. See for example Refs. [8, 56] where candid discussions are presented on the limitations of matching decoder variations correcting for depolarising noise. In this work, we introduce a syndrome lattice for the surface code that accounts for correlations between bit-flip and dephasing errors, which we obtain by exploiting a correspondence between the surface code and the color code, commonly known as an unfolding map [58, 59, 60, 61]. Through this mapping, we can adopt color code decoding algorithms to correct noise on the surface code. Indeed, a syndrome lattice for our decoder naturally ensues by adapting the decoding methods introduced in Ref. [48].
@@
We find that our unified decoder determines a least-weight correction for all weight $d/2$ depolarising noise errors for surface codes with even distance $d$. As such, in the limit of low error rate, we obtain optimal logical failure rates. In contrast, conventional matching decoders that do not account for correlated errors have a logical failure rate $O(2^{d/2})$ a factor higher than the decoder we present. We adopt a number of methods to analyse the performance of our decoder at low error rates. We first evaluate the number of weight $d/2$ errors that should lead to a logical failure for both our matching decoder, as well as a conventional matching decoder. We verify our expressions with exhaustive searching of weight $d/2$ errors. In addition, we measure the logical failure rate of our decoder using the splitting method [62, 63]. We find the numerical results we obtain to be consistent with our analytic results in the limit of vanishing physical error rate, thereby verifying the performance of our decoder in comparison to conventional decoders. We also show that the threshold of our decoder is comparable to more conventional decoding methods. Finally, we investigate the performance of our scheme to decode the color code undergoing bit-flip noise. We find our analysis sheds light on the role of entropy in quantum error correction with the matching decoders that are central to our study [64, 65, 66].
@@ 2.1 The stabilizer formalism
Quantum-error correcting codes are designed to protect logical information that is encoded in a subspace of a larger physical Hilbert space. This subspace is called the code space. A large class of QEC codes are readily described using the stabilizer formalism [67], whereby a QEC code is defined by an Abelian subgroup of the Pauli group $\mathcal{P}$, known as the stabilizer group $\mathcal{S}\subset\mathcal{P}$. Up to phases, the Pauli group on $n$ qubits is generated by the operators $n$-qubit Pauli matrices $X_{q}$ and $Z_{q}$ which respectively denote the standard Pauli-X and Pauli-Z operatorss acting on qubit $q$. The code space of a stabilizer code is the +$1$ eigenspace of all of its elements, namely $s\ket{\psi}=\ket{\psi}$ $\forall$ $s\in\mathcal{S}$, where the code space is spanned by eigenstates |$\ket{\psi}$.
@@
A correction $C$ can be found and applied to return the code to the ground subspace such that $CE\in\mathcal{S}$, and recover the encoded logical state |$\ket{\psi}$. We obtain $C$ using a decoding algorithm that is designed to determine a highly probable error giving rise to a measured syndrome, given a known error model where we assume errors are introduced at a low rate. In general, an $\llbracket n,k,d\rrbracket$ quantum code is characterised by the number of physical qubits $n$, the number of logical qubits $k$ that they encode, and a code distance $d$ given by the minimum weight of an error configuration triggering a logical error.
@@ 2.2 The surface code
In this work, we propose and implement a matching decoder to correct for the surface code undergoing an independent and identically distributed depolarising noise model. This error model is such that each qubit experiences either a Pauli-$X$, -$Y$, or -$Z$ error with probability $p/3$, or no error with probability $(1-p)$. We indicate examples of these single-qubit errors and their corresponding syndromes in Fig. 1.
@@ 2.3 The color code
The color code is defined on a three-colorable two-dimensional lattice. We will focus on the regular square-octagon color code with square boundaries, see Fig. 2. Physical qubits are located on the vertices of the lattice, $v$, and each face $f$ supports two stabilizer operators $S_{f}^{P}=\prod_{v\in E_{f}}P_{e}$ where $P\in\{X,Z\}$ and $E_{f}$ denotes the qubits on the boundaries of face $f$.
We define a set of color labels red, green and blue in bold $\mathcal{C}=\{\textbf{r},\textbf{g},\textbf{b}\}$, and each face of the code is assigned a label $\textbf{u}\in\mathcal{C}$ such that no two faces of the same color touch.
@@ 3 Materialised symmetries and decoding
In this section, we review the concept of materialised symmetries for topological quantum codes introduced in Refs. [15, 41, 48, 55]. These symmetries give rise to a parity conservation law on syndrome defects, that is, any error creates an even number of defects on the symmetry. We exploit this parity conservation law to make a matching decoder for the square-boundary color code. In Ref. [48], a global symmetry termed the unified lattice was derived for the triangular color code with a single red, blue and green boundary. Notably, the unified lattice presented in Ref. [48] had the topology of a Möbius strip. Due to the unified lattice’s improved capacity to handle correctable errors at the boundary [68], Ref. [48] reported improvements in both the threshold and logical failure rates for the triangular color code using decoding on the unified lattice. Motivated by this, we introduce a unified decoder for the color code presented in Sec. 2.3. In addition to introducing the unified lattice decoder, we will also review the matching decoder on the restricted lattice [44, 45, 46, 47, 49]. In later sections, we will compare the restricted-lattice decoder to the unified-lattice decoder.
@@ 3.1 Symmetries of the color code
We find the symmetries of the color code by considering subsets of its stabilizer generators with specific colors [45, 48]. Our goal is to derive a single ‘unifying’ symmetry for the color code with boundaries. We also investigate other color code decoders based on restricted lattices. Interestingly, we find that the square-boundary color code has a unified lattice with a topology that is distinct from that in Ref. [48], and is given by the manifold in Fig. 3. In this subsection, we show how to derive this unified lattice for the square boundary color code, which we will adopt in our matching decoder implementation.
@@
We show the restricted lattices $\mathcal{R}_{\mathbf{b}},\mathcal{R}_{\mathbf{r}},$ and $\mathcal{R}_{\mathbf{g}}$ in Fig. 4(a). Errors in the bulk of the color code preserve defect parity on any single restricted lattice. We show examples of bulk errors on the restricted lattices in Fig. 4(a), where the bottom left error on the lattice flips a single qubit and creates a single defect of each color, and the top right error flips two qubits, and creates two green defects.
@@
where we take the product over all faces of the restricted lattice, i.e. those not colored $\mathbf{u}$. This operator is highlighted for each restricted lattice in Fig. 4(b). We note that the boundary operator $b_{\mathbf{u}}$ is supported on all boundaries and corners that are not $\mathbf{u}$-colored.
@@
3.2 Matching decoders and code symmetries
@@
The unified lattice is obtained from the global symmetry of the color code described in the previous subsection, on which errors give rise to pairs of defects. Given a collection of defects on the unified lattice, the matching-based decoding problem reduces to finding a pairing between these defects that is created by an underlying error with maximal probability. Towards this goal, a matching between a pair of defects is assigned a weight inversely dependent on the probability of errors on all the qubits along the shortest path connecting the two. As a result, decoding is equivalent to finding the perfect matching of minimum weight on this weighted graph on the unified lattice.
@@
Given a matching on a particular lattice – unified or restricted – we now lay out the procedure to obtain the relevant correction to the logical qubit subjected to $X$ errors on the underlying physical qubits. Note that the following procedure is valid for both the restricted and unified lattices. We first apply the Pauli-$Z$ corrections to qubits along the shortest paths determined by the matching algorithm. This returns the logical qubit to the code space. During this process, we track the parity of a logical operator, here the vertical $Z$ logical depicted in Fig. 5. We successfully return the encoded qubit to the correct logical state if the decoding algorithm has determined that the parity of matched paths across this logical is the same as the parity of physical Pauli errors on the qubits along the logical, a method discussed in more detail in Ref. [55, 48]. For example, in Fig. 5, if the decoder finds the yellow matching, the corresponding correction, together with the initial error, applies $\overline{X}$ to the encoded qubit; the grey matching returns the lattice to the correct state. Henceforth, we will refer to the MWPM-based decoder applied to syndrome data on the restricted (unified) lattice as the restricted (unified) decoder.
4 Mapping surface code errors to color code errors
@@
We have introduced the surface code and the color code in the previous section, and proposed a decoding procedure for bit-flip errors acting on the color code by matching on either the restricted lattice or the unified lattice. We now describe the local mapping from the surface code onto the square-octagonal color code, inspired by the unfolding of the color code [59], and introduce a mapping from depolarising errors on the former to correlated bit-flip errors on the latter. We exploit this mapping to adopt the restricted-lattice and unified-lattice color code decoder for the surface code undergoing depolarising noise.
@@
Let us begin by stating the equivalence between the different lattice elements under this mapping, by considering the surface code displayed with a blue lattice in Fig. 6(a). First, we map the vertex and face operators of this blue surface code onto octagonal face operators of the color code. Specifically, each surface code star operator corresponds to a blue octagon $S_{f}$ of the color code and, likewise, each face operator of the surface code corresponds to a green octagonal face operator in the color code picture $S_{f}$. This is shown in Fig. 6(a), where the star and face operators of the blue surface code align with their corresponding color code stabilizers exactly, thereby illustrating this mapping. We also show the surface code displayed with a green lattice, which represents a rotation of the blue counterpart, and where a similar argument holds.
Now that we have identified stars and plaquettes of each surface code with octagonal face operators of the color code, we must choose errors in the color code picture that give rise to equivalent syndromes in the surface code picture. A Pauli-$Z$ error on the surface code will violate its two adjacent vertex operators, as shown in Fig. 1. In the color code picture, this error will violate the corresponding blue octagonal operators as shown in Fig. 6(b)(i) where we indicate this $Z$-flip on the blue surface code edge. To obtain this violation in the color code picture, we add two horizontal bit-flip errors onto the single red face of the color code on which the edge lies such that its two adjacent blue octagons support a syndrome, see Fig. 6(a). Similarly, a Pauli-$X$ error in the surface code picture will violate two adjacent face operators, which correspond to green octagons in the color code picture in Fig. 6(b)(ii). We find that the error in the color code picture that violates these appropriate color code octagons is a pair of bit-flip errors lying vertically on a single red square-shaped face operator shown in the corresponding diagram in Fig. 6(a).
@@
We have now proposed an error model that violates pairs of either blue or green face operators according to the mapping from the surface code undergoing a depolarising noise model. As we have alluded, single qubit errors in the surface code picture correspond to two-qubit errors supported on individual red faces of the color code. Indeed, we see that the edges in the surface code picture each have an underlying red face operator in the color code picture, where the red square inherits an orientation from the underlying edge. We therefore identify qubits in the surface code picture with red faces in the color code picture. In what follows, it will be helpful to keep this identification in mind as we describe the mapping more rigorously. We note that this mapping is non-trivial in the sense that the commutation relations of the errors are not preserved. Nevertheless, we find a valid decoder for the surface code with the equivalence we have proposed.
@@
We then express the stabilizers of the blue and green surface codes respectively, using the two logical operators of the inner code $\{\overline{X}_{1},\overline{Z}_{1}\}$ and $\{\overline{X}_{2},\overline{Z}_{2}\}$, where the first(second) qubit encodes the blue(green) copy of the surface code, as shown in Fig. 6(d) (Fig. 6(e)). We emphasise that that the orientation of the edge qubit of the surface code dictates the orientation of the encoding of the $\llbracket 4,2,2\rrbracket$ code. See Fig. 6(c). Indeed, a red square associated with a horizontal (vertical) edge of the blue (green) surface code is encoded as per Fig. 6(c), but when traversed by a vertical edge, this encoding must be rotated by 90° to preserve our mapping.
@@
Finally, since the surface code on the green lattice is a duplicate of its blue counterpart up to a Hadamard rotation, we identify certain logical operators of the inner code to complete our mapping. Indeed, under the duplication, with the Hadamard rotation, we have the following relations between pairs of logical operators of the inner code, namely $\overline{X}_{1}\equiv\overline{Z}_{2}$ and $\overline{Z}_{1}\equiv\overline{X}_{2}$. This ensures the syndrome correspondence in Fig. 6(a) such that all depolarising errors on the surface code now have a one-to-one correspondence with a commuting two-qubit bit-flip error model. Most importantly, this enables us to use a decoder for a CSS code [3, 73] to correct after depolarising noise on the surface code, represented in the color code picture.